При каких значениях а и b уравнение ax+5=7x+b не имеет решения
Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

can u write this in english

Answer by Alan3354(69443)

For what values ​​of a and b of the equation ax +5 = 7x + b has no solution
--------------------
x*(a-7) = b-5
x = (b-5)/(a-7)
@ a = 7 --> undefined